OpenStudy (anonymous):

Simplify. Write your answers without exponents. (1/32)^-3/5 AND 81^-3/4

OpenStudy (anonymous):

\[\huge x ^{\frac{ m }{ n }}=\sqrt[n]{x^m} = (\sqrt[n]{x})^m\]

OpenStudy (anonymous):

yayy! thanx!

OpenStudy (anonymous):

Np :)